OpenStudy (anonymous):

Find the product of (x − 3)^2 x^2-9?

OpenStudy (anonymous):

\[(x - 3)^2 \times (x^2 -9)\] ?

OpenStudy (anonymous):

the question is the first line i just wanna know if my answer is correct on the second line?

OpenStudy (anonymous):

\[(x -3)^3 \times (x +3)\] Note. If the question is the one I supposed

OpenStudy (anonymous):

yes but not the exponent of 3

OpenStudy (anonymous):

\[x^2 -9 = (x-3)(x+3)\]

OpenStudy (anonymous):

then I get \[(x-3)(x+3)^3\]

OpenStudy (anonymous):

So your question is \[(x-3)^2=(x-3)(x-3)\] this can be broken down further so you can see where I am getting the answer from of \[(x-3)(x-3)=x \times (x-3)-3 \times (x-3)\] taking the first terms on the RHS we have \[x(x-3)=x \times x +x \times (-3) = x^2 -3x \] and \[-3(x-3) = -3 \times x + (-3) \times (-3)=-3x +9\] Putting this together then you have \[(x-3)(x-3)=x^2-3x-3x+9=x^2-6x+9\]